# Quantum metrics from length functions on quantum groups

### v1

@@ 1. Introduction
Building on the classical Gelfand correspondence between compact Hausdorff spaces and commutative unital $C^{*}$-algebras, numerous non-commutative analogues of classical point-set based theories have emerged over the past 50 years. Prominent examples include the theory of quantum groups [KS97], Connes’ noncommutative geometry [Con94] and Rieffel’s theory of compact quantum metric spaces [Rie04a], the latter of which is the primary object of study in the present paper. As the name suggests, Rieffel’s theory is a non-commutative extension of the theory of compact metric spaces, and has provided a mathematically rigorous framework within which heuristic statements from physics can be proven rigorously. A prime example of this phenomenon is Rieffel’s seminal result that matrix algebras (fuzzy spheres) converge to the 2-sphere [Rie04c].
More formally, a compact quantum metric space consists of an operator system ${\mathcal{X}}$ equipped with a seminorm $L\colon{\mathcal{X}}\to[0,\infty)$ such that the associated Monge-Kantorovič metric
@@
metrises the weak topology on the state space ${\mathcal{S}}({\mathcal{X}})$; see section 2.1 for the precise requirements and examples.
The theory of compact quantum metric spaces is closely linked to Connes’ non-commutative (differential) geometry by means of the following construction: Given a spectral triple $({\mathcal{A}},H,D)$, one obtains a seminorm $L_{D}$ on ${\mathcal{A}}$ by setting $L_{D}(a):=\|[D,a]\|$, and this construction often (but not always) yields interesting examples of compact quantum metric spaces [OR05, CR17, AK18, CI06]. It is generally a non-trivial task to verify if a given spectral triple gives rise to a compact quantum metric space, and even some of the most fundamental examples in the theory are not fully understood. As an example, if $\Gamma$ is a discrete group equipped with a proper length function $\ell\colon\Gamma\to[0,\infty)$, one obtains a natural spectral triple $({\mathbb{C}}\Gamma,\ell^{2}(\Gamma),D_{\ell})$ (see Section 2.1 for details) and the only classes of groups for which it is known that $L_{\ell}:=L_{D_{\ell}}$ provides a compact quantum metric structure, are hyperbolic groups [OR05] and groups of polynomial growth [CR17]. In both cases, the argument is rather involved, and except for minor improvements [LW21], no new examples have been obtained since Christ and Rieffel’s paper [CR17] from 2017. It therefore seems natural to attempt to expand the number of examples by allowing more flexibility in the defining data, and the present paper implements this strategy by extending the domain from groups to quantum groups.
@@
is able to detect when $(C({\mathbb{G}}),L_{\ell})$ is a compact quantum metric space. Here $\sigma$ denotes the flip-automorphism on $C({\mathbb{G}})\otimes C({\mathbb{G}})$.
@@
If ${\mathbb{G}}$ is a compact, coamenable quantum group of Kac type and $\ell\colon\operatorname{Irred}({\mathbb{G}})\to[0,\infty)$ is a proper length function, then $(C({\mathbb{G}}),L_{\ell})$ is a compact quantum metric space if and only if $(C_{z}({\mathbb{G}}),L_{\ell})$ is a compact quantum metric space.
Here the notion of coamenability is the quantum analogue of amenability of a discrete group, while the Kac type assumption can be viewed as unimodularity of the discrete dual quantum group — an assumption that is of course vacuous in the setting of discrete group where unimodularity is automatic; see Section 2.2 for more details.
@@
Let ${\mathbb{G}}_{1}$ and ${\mathbb{G}}_{2}$ be compact quantum groups with length functions $\ell_{1}$ and $\ell_{2}$ and assume that $\ell_{1}$ defines a compact quantum metric structure on $F({\mathbb{G}}_{1})$. If there exists a bijection $\alpha\colon\operatorname{Irred}({\mathbb{G}}_{1})\to\operatorname{Irred}({\mathbb{G}}_{2})$ which intertwines the length functions and extends to a *-isomorphism $F({\mathbb{G}}_{1})\simeq F({\mathbb{G}}_{2})$ and if ${\mathbb{G}}_{2}$ is coamenable and of Kac type, then $\ell_{2}$ yields a compact quantum metric structure on both $F({\mathbb{G}}_{2})$ and $C({\mathbb{G}}_{2})$.
The quantum counterpart to a classical compact Lie group is known as a compact matrix quantum group, and the additional data is encoded in a fixed (not necessarily irreducible) corepresentation $u$ whose matrix coefficients generate $\operatorname{Pol}({\mathbb{G}})$ as a *-algebra. One refers to the chosen $u$ as the fundamental corepresentation. Natural examples of matrix quantum groups include $SU_{q}(2)$ as well as Wang’s liberated versions of the permutation groups, orthogonal groups and unitary groups, denoted $S_{n}^{+},O_{n}^{+}$ and $U_{n}^{+}$, respectively. Note that $SU_{q}(2)$ is coamenable for all $q$ but only of Kac type when $q=\pm 1$, while the liberated quantum groups are all of Kac type, but only $O_{2}^{+},S_{2}^{+},S_{3}^{+}$ and $S_{4}^{+}$ are coamenable. In the context of the present paper, the most interesting quantum permutation group is $S_{4}^{+}$, since $C(S_{4}^{+})$ is an infinite-dimensional, non-commutative $C^{*}$-algebra while $C(S_{n}^{+})=C(S_{n})$ for $n\leqslant 3$, so that there are no non-classical quantum symmetries. Note also that one has $SU_{-1}(2)\simeq O_{2}^{+}$ and that $S_{4}^{+}\simeq SO_{-1}(q)$ (a deformation of the classical Lie group $SO(3)$); see [BB09, Ban99b, Wan98] for more details. All of these comments are meant to provide insight into why these are the examples that appear in Corollary C below.
When ${\mathbb{G}}$ is a matrix quantum group and the fundamental corepresentation $u$ is equivalent to its conjugate corepresentation, one has that every $\alpha\in\operatorname{Irred}({\mathbb{G}})$ appears in some tensor power $u^{\boxtimes k}$, and ${\mathbb{G}}$ thereby obtains a natural length function $\ell\colon\operatorname{Irred}({\mathbb{G}})\to{\mathbb{N}}_{0}$ by setting $\ell(\alpha)$ equal to the smallest such $k$. This is the situation when ${\mathbb{G}}$ is $SU(2)$, $O_{2}^{+}$, $SO(3)$ or $S_{4}^{+}$ and for these examples we obtain quantum metric spaces:
@@
Throughout the article, we will assume that inner products are conjugate linear in the second argument. Moreover, $\odot$ will denote the algebraic tensor product of algebras, $\bar{\otimes}$ is the von Neumann tensor product, and $\hat{\otimes}$ is the (completed) tensor product of Hilbert spaces. We will denote by $\otimes$ the minimal tensor product of $C^{*}$-algebras or the algebraic tensor product of vector spaces, and the usage will be clear from context. For indices $i$ and $j$ in some index set, we will denote by $\delta_{ij}$ the corresponding Kronecker delta.
@@ Acknowledgements
The authors gratefully acknowledge the financial support from the Independent Research Fund Denmark through grant no. 9040-00107B, 7014-00145B and 1026-00371B and from the the European Commision through the MSCA Staff Exchanges grant no. 101086394. The first named author was supported by The Research Council of Norway project 324944.
@@ 2.1. Compact quantum metric spaces
The present section contains a brief introduction to Rieffel’s theory of compact quantum metric spaces [Rie99, Rie04b, Rie04a], which provides an elegant non-commutative analogue of the classical theory of compact metric spaces. The theory of compact quantum metric spaces can be based on either order unit spaces [Rie99], $C^{*}$-algebras [Li09, Lat16] or operator systems [Ker03, KL09], among which we have chosen the latter approach for the sake of coherence with recent developments in [KK25, van21, Cv21]. To this end, we recall that a (concrete) operator system is a self-adjoint, unital subspace ${\mathcal{X}}$ in a given unital $C^{*}$-algebra $A$. We will denote the closure of ${\mathcal{X}}$ by $X$ and refer to ${\mathcal{X}}$ as complete if ${\mathcal{X}}=X$. An element $x\in{\mathcal{X}}$ is positive if this is the case in the ambient $C^{*}$-algebra $A$, and the state space ${\mathcal{S}}({\mathcal{X}})$ is defined as the positive, unital functionals on ${\mathcal{X}}$. Note that ${\mathcal{S}}({\mathcal{X}})$ is compact for the weak topology and homeomorphic to ${\mathcal{S}}(X)$ via the restriction map.
Definition 2.1 ([Rie16]).
A slip-norm on an operator system ${\mathcal{X}}$ is a seminorm $L\colon{\mathcal{X}}\to[0,\infty)$ satisfying that $L(1)=0$ and $L(x)=L(x^{*})$ for all $x\in{\mathcal{X}}$.
@@
A compact quantum metric space consists of a pair $({\mathcal{X}},L)$ where ${\mathcal{X}}$ is an operator system and $L\colon{\mathcal{X}}\to[0,\infty)$ is a slip-norm such that $d_{L}$ metrises the weak topology on ${\mathcal{S}}({\mathcal{X}})$. In this case, $L$ is referred to as a Lip-norm.
@@
${\mathcal{X}}:=C_{\text{Lip}}(M):=\{f\in C(M)\mid f\text{ is Lipschitz continuous }\},$
and defines $L_{d}\colon{\mathcal{X}}\to[0,\infty)$ by setting
@@
[…]
$L(a):=\sup\left\{\frac{\|\alpha_{g}(a)-a\|}{d_{G}(g,a)}\mid g\in G\setminus\{e\}\right\}$
on those elements for which the right hand side is finite [Rie98, Theorem 2.3].
@@
As is customary in the field, we will often extend a slip-norm $L$ on an operator system ${\mathcal{X}}$ to the closure $X$ by declaring $L(x)=\infty$ for $x\in X\setminus{\mathcal{X}}$. We will further employ the standard abuse of notation and also refer to $(X,L)$ as a compact quantum metric space when $L\colon{\mathcal{X}}\to[0,\infty)$ is a Lip-norm.
The theory of compact quantum metric spaces is heavily inspired by Connes’ non-commutative differential geometry [Con94], in that many interesting examples arise from spectral triples. For our purposes, it is most convenient to choose a minimal set of axioms defining a spectral triple, and we will thus be ignoring gradings, regularity properties etc. For background on non-commutative geometry, we refer to [Con94, GBVF01].
@@
    Every $a\in{\mathcal{A}}$ preserves the domain of $D$ and $[D,a]=Da-aD$ extends to a bounded operator $\partial_{D}(a)$ on $H$.
@@
Given a spectral triple $({\mathcal{A}},H,D)$, one obtains a natural seminorm on ${\mathcal{A}}$ by setting $L_{D}(a):=\|\partial_{D}(a)\|$. In many interesting cases [AK18, Rie02, CI06], but not always [KN24, JP16], $L_{D}$ is a a Lip-norm in which case $({\mathcal{A}},H,D)$ is called a spectral metric space.
To illustrate the complexity of the situation, we now consider the fundamental example, dating back to Connes’ seminal paper [Con89], arising from a countable discrete group $\Gamma$ equipped with a proper length function $\ell\colon\Gamma\to[0,\infty)$. That is, the function $\ell$ satisfies
@@
From a length function $\ell$, one obtains a spectral triple $({\mathbb{C}}\Gamma,\ell^{2}(\Gamma),D_{\ell})$ on the reduced group $C^{*}$-algebra $C^{*}_{{\operatorname{red}}}(\Gamma)$, by defining $D_{\ell}$ to be the closure of the unbounded operator $\mathcal{D}_{\ell}\colon\text{span}\{\delta_{x}\mid x\in\Gamma\}\to\ell^{2}(\Gamma)$ defined on the standard orthonormal basis by $\mathcal{D}_{\ell}(\delta_{x}):=\ell(x)\cdot\delta_{x}$. Even in this case, it is very unclear when $L_{D_{\ell}}$ is a Lip-norm — at the time of writing, this is basically only known for groups of polynomial growth [CR17] and hyperbolic groups [OR05].
@@
Theorem 2.6 ([Rie98, Theorem 1.8]).
The pair $({\mathcal{X}},L)$ is a compact quantum metric space if and only if the set $\{x\in{\mathcal{X}}\mid L(x)\leqslant 1\}$ projects to a totally bounded set in ${\mathcal{X}}/{\mathbb{C}}$ via the quotient map ${\mathcal{X}}\ni x\mapsto[x]\in{\mathcal{X}}/{\mathbb{C}}$.
@@
$\|[x]\|_{{\mathcal{X}}/{\mathbb{C}}}\leqslant C\cdot L(x)\quad\mbox{for all }x\in\mathcal{X}.$ (2.1)
Having finite diameter, in the sense just defined, is equivalent to the (extended) metric $d_{L}$ assigning a finite diameter to the state space ${\mathcal{S}}({\mathcal{X}})$. More precisely, the smallest constant $C$ satisfying the inequality (2.1) will be equal to half the diameter of the (extended) metric space $({\mathcal{S}}({\mathcal{X}}),d_{L})$; [Rie99, Proposition 2.2]. Note also, that by connectedness and compactness of ${\mathcal{S}}(\mathcal{X})$ in the weak topology, a compact quantum metric space automatically has finite diameter. Kaad’s criterion now reads as follows:
Theorem 2.8 ([Kaa24, Theorem 3.1]).
@@
    For every ${\varepsilon}>0$ there exist an operator system ${\mathcal{X}}_{\varepsilon}$ and unital, bounded maps $\iota,\Phi\colon{\mathcal{X}}\to{\mathcal{X}}_{\varepsilon}$ such that $\iota$ is an isometry, $\Phi$ is positive with finite-dimensional image and $\|\iota(x)-\Phi(x)\|\leqslant{\varepsilon}\cdot L(x)$ for all $x\in{\mathcal{X}}$.

### v3

@@ 1. Introduction
Building on the classical Gelfand correspondence between compact Hausdorff spaces and commutative unital $C^{*}$-algebras, numerous non-commutative analogues of classical point set based theories have emerged over the past 50 years. Prominent examples include the theory of quantum groups [KlSc:QGR], Connes’ noncommutative geometry [Con:NCG] and Rieffel’s theory of compact quantum metric spaces [Rie:CQM], the latter of which is the primary object of study in the present paper. As the name suggests, Rieffel’s theory is a non-commutative extension of the theory of compact metric spaces, and has provided a mathematically rigorous framework within which heuristic statements from physics can be proven rigorously. A prime example of this phenomenon is Rieffel’s seminal result that matrix algebras (fuzzy spheres) converge to the 2-sphere [Rie:MSG].
More formally, a compact quantum metric space consists of an operator system $\mathcal{X}$} equipped with a seminorm $L\colon{\mathcal{X}}\rightarrow[0,\infty)$ such that the associated Monge-Kantorovič metric
@@
metrizes the weak topology on the state space ${\mathcal{S}}({\mathcal{X}})$; see Section 2.1 for the precise requirements and examples.
The theory of compact quantum metric spaces is closely linked to Connes’ non-commutative (differential) geometry by means of the following construction: Given a spectral triple $({\mathcal{A}},H,D)$, one obtains a seminorm $L_{D}$ on $\mathcal{A}$} by setting $L_{D}(a):=\|[D,a]\|$, and this construction often (but not always) yields interesting examples of compact quantum metric spaces [OzawaRieffel2005, ChristRieffel2017, AgKa:PSM, Christensen-Ivan:spectral-triples]. It is generally a non-trivial task to verify if a given spectral triple gives rise to a compact quantum metric space, and even some of the most fundamental examples in the theory are not fully understood. As an example, if $\Gamma$ is a discrete group equipped with a proper length function $\ell\colon\Gamma\rightarrow[0,\infty)$, one obtains a natural spectral triple $({\mathbb{C}}\Gamma,\ell^{2}(\Gamma),D_{\ell})$ (see Section 2.1 for details) and the only classes of groups for which it is known that $L_{\ell}:=L_{D_{\ell}}$ provides a compact quantum metric structure, are hyperbolic groups [OzawaRieffel2005] and groups of polynomial growth [ChristRieffel2017]. In both cases, the argument is rather involved, and except for minor improvements [LongWu:twisted], no new examples have been obtained since Christ and Rieffel’s paper [ChristRieffel2017] from 2017. It therefore seems natural to attempt to expand the number of examples by allowing for more flexibility in the defining data, as was done for instance in [GerontogiannisMesland2025] and [Antonescu-Christensen], and the present paper suggests a new direction to this strategy by extending the domain from groups to quantum groups.
@@
is able to detect when $(C({\mathbb{G}}),L_{\ell})$ is a compact quantum metric space. Here $\sigma$ denotes the flip automorphism on $C({\mathbb{G}})\otimes C({\mathbb{G}})$.
@@
If $\mathbb{G}$} is a compact, coamenable quantum group of Kac type and $\ell\colon\operatorname{Irred}({\mathbb{G}})\rightarrow[0,\infty)$ is a proper length function, then $(C({\mathbb{G}}),L_{\ell})$ is a compact quantum metric space if and only if $(C_{z}({\mathbb{G}}),L_{\ell})$ is a compact quantum metric space.
Here the notion of coamenability is the quantum analogue of amenability of a discrete group, while the Kac type assumption can be viewed as unimodularity of the discrete dual quantum group — an assumption that is of course vacuous in the setting of discrete groups where unimodularity is automatic; see Section 2.2 for more details.
The proof of Theorem A builds on the approximation techniques developed in [AKK:Podcon], the main novelty being the utilization of a sequence of central functionals approximating the counit. The existence of such a sequence is ensured by the coamenability assumption, while the Kac type assumption provides a conditional expectation $E\colon C({\mathbb{G}})\rightarrow C_{z}({\mathbb{G}})$ which, in turn, ensures that the Monge-Kantorovič distance between two central states can be detected by their restriction to the algebra of central functions (see Lemma 5.3).
@@
Let ${\mathbb{G}}_{1}$ and ${\mathbb{G}}_{2}$ be compact quantum groups with length functions $\ell_{1}$ and $\ell_{2}$ and assume that $\ell_{1}$ defines a compact quantum metric structure on $F({\mathbb{G}}_{1})$. If there exists a bijection $\alpha\colon\operatorname{Irred}({\mathbb{G}}_{1})\rightarrow\operatorname{Irred}({\mathbb{G}}_{2})$ which intertwines the length functions and extends to a *-isomorphism $F({\mathbb{G}}_{1})\simeq F({\mathbb{G}}_{2})$, and if ${\mathbb{G}}_{2}$ is coamenable and of Kac type, then $\ell_{2}$ yields a compact quantum metric structure on both $F({\mathbb{G}}_{2})$ and $C({\mathbb{G}}_{2})$.
The quantum counterpart to a classical compact Lie group is known as a compact matrix quantum group, and the additional data is encoded in a fixed (not necessarily irreducible) corepresentation $u$ whose matrix coefficients generate $\operatorname{Pol}({\mathbb{G}})$ as a *-algebra. One refers to the chosen $u$ as the fundamental corepresentation. Natural examples of matrix quantum groups include $SU_{q}(2)$ as well as Wang’s liberated versions of the permutation groups, orthogonal groups and unitary groups, denoted $S_{n}^{+},O_{n}^{+}$ and $U_{n}^{+}$, respectively. Note that $SU_{q}(2)$ is coamenable for all $q$ but only of Kac type when $q=\pm 1$, while the liberated quantum groups are all of Kac type, but only $O_{2}^{+},S_{2}^{+},S_{3}^{+}$ and $S_{4}^{+}$ are coamenable. In the context of the present paper, the most interesting quantum permutation group is $S_{4}^{+}$, since $C(S_{4}^{+})$ is an infinite-dimensional, non-commutative $C^{*}$-algebra while $C(S_{n}^{+})=C(S_{n})$ for $n\leqslant 3$, so that there are no non-classical quantum symmetries. Note also that one has $SU_{-1}(2)\simeq O_{2}^{+}$ and that $S_{4}^{+}\simeq SO_{-1}(3)$ (a deformation of the classical Lie group $SO(3)$); see [banica-bichon, banica-sym, Wang:quantum-symmetry] for more details. All of these comments are meant to provide insight into why these are the examples that appear in Corollary C below.
When $\mathbb{G}$} is a matrix quantum group and the fundamental corepresentation $u$ is equivalent to its conjugate corepresentation, one has that every $\alpha\in\operatorname{Irred}({\mathbb{G}})$ appears in some tensor power $u^{\boxtimes k}$, and $\mathbb{G}$} thereby obtains a natural length function $\ell\colon\operatorname{Irred}({\mathbb{G}})\rightarrow{\mathbb{N}}_{0}$ by setting $\ell(\alpha)$ equal to the smallest such $k$. This is the situation when $\mathbb{G}$} is $SU(2)$, $O_{2}^{+}$, $SO(3)$ or $S_{4}^{+}$, and for these examples we obtain quantum metric spaces:
@@
Throughout the article, we will assume that inner products are conjugate linear in the first argument. Moreover, $\odot$ will denote the algebraic tensor product of algebras, $\bar{\otimes}$ is the von Neumann tensor product, and $\hat{\otimes}$ is the (completed) tensor product of Hilbert spaces. We will denote by $\otimes$ the minimal tensor product of $C^{*}$-algebras or the algebraic tensor product of vector spaces, and the usage will be clear from context. For indices $i$ and $j$ in some index set, we will denote by $\delta_{ij}$ the corresponding Kronecker delta.
@@ Acknowledgements
The authors would like to thank Jens Kaad for his valuable insights and many interesting discussions, Christian Voigt for reminding them about the importance of understanding basic examples and Dimitris M. Gerontogiannis and Bram Mesland for asking interesting questions and pointing out the new approach in [GerontogiannisMesland2025]. Lastly, they would like to thank the anonymous referee for their thorough reading of the paper and many helpful comments.
The authors gratefully acknowledge the financial support from the Independent Research Fund Denmark through grant no. 9040-00107B, 7014-00145B and 1026-00371B and from the European Commision through the MSCA Staff Exchanges grant no. 101086394. The first named author was supported by The Research Council of Norway project 324944.
@@ 2.1. Compact quantum metric spaces
The present section contains a brief introduction to Rieffel’s theory of compact quantum metric spaces [Rie:MSS, Rie:GHD, Rie:CQM], which provides an elegant non-commutative analogue of the classical theory of compact metric spaces. The theory of compact quantum metric spaces can be based on either order unit spaces [Rie:MSS], $C^{*}$-algebras [Li:ECQ, Lat:QGH] or operator systems [Ker:MQG, Kerr-Li:GH], among which we have chosen the latter approach for the sake of coherence with recent developments in [KaadKyed2022, walter:GH-convergence, walter-connes:truncations]. To this end, we recall that a (concrete) operator system is a self-adjoint, unital subspace $\mathcal{X}$} in a given unital $C^{*}$-algebra $A$. We will denote the closure of $\mathcal{X}$} by $X$ and refer to $\mathcal{X}$} as complete if ${\mathcal{X}}=X$. An element $x\in{\mathcal{X}}$ is positive if this is the case in the ambient $C^{*}$-algebra $A$, and the state space ${\mathcal{S}}({\mathcal{X}})$ is defined as the positive, unital functionals on $\mathcal{X}$}. Note that ${\mathcal{S}}({\mathcal{X}})$ is compact for the weak topology and homeomorphic to ${\mathcal{S}}(X)$ via the restriction map.
Definition 2.1 ([Rie:matricial-bridges]).
A slip-norm on an operator system $\mathcal{X}$} is a seminorm $L\colon{\mathcal{X}}\rightarrow[0,\infty)$ satisfying that $L(1)=0$ and $L(x)=L(x^{*})$ for all $x\in{\mathcal{X}}$.
@@
A compact quantum metric space consists of a pair $({\mathcal{X}},L)$ where $\mathcal{X}$} is an operator system and $L\colon{\mathcal{X}}\rightarrow[0,\infty)$ is a slip-norm such that $d_{L}$ metrizes the weak topology on ${\mathcal{S}}({\mathcal{X}})$. In this case, $L$ is referred to as a Lip-norm.
@@
[…]
and defines $L_{d}\colon{\mathcal{X}}\rightarrow[0,\infty)$ by setting
@@
In this way, one obtains a compact quantum metric space. This was originally proven in [KaRu:FSE, KaRu:OSC]; for a modern approach see [Kaad2023, Lemma 3.4]. Interesting non-commutative examples are obtained by considering a strongly continuous, ergodic action $G\overset{\alpha}{\curvearrowright}A$ of a compact metric group $G$ on a unital $C^{*}$-algebra $A$ and defining
$L(a):=\sup\left\{\frac{\|\alpha_{g}(a)-a\|}{d_{G}(g,e)}\mid g\in G\setminus\{e\}\right\}$
on those elements for which the right hand side is finite [Rie:MSA, Theorem 2.3].
@@
As is customary in the field, we will often extend a slip-norm $L$ on an operator system $\mathcal{X}$} to the closure $X$ by declaring $L(x)=\infty$ for $x\in X\setminus{\mathcal{X}}$. We will further employ the standard abuse of notation and also refer to $(X,L)$ as a compact quantum metric space when $L\colon{\mathcal{X}}\rightarrow[0,\infty)$ is a Lip-norm.
The theory of compact quantum metric spaces is heavily inspired by Connes’ non-commutative differential geometry [Con:NCG], in that many interesting examples arise from spectral triples. For our purposes, it is most convenient to choose a minimal set of axioms defining a spectral triple, and we will thus be ignoring gradings, regularity properties etc. For background on non-commutative geometry, we refer to [Con:NCG, elements-of-ncg].
@@
    Every $a\in{\mathcal{A}}$ preserves the domain of $D$ and [D,a]=DaaD[D,a]=Da-aD extends to a bounded operator $\partial_{D}(a)$ on $H$.
@@
Given a spectral triple $({\mathcal{A}},H,D)$, one obtains a natural seminorm on $\mathcal{A}$} by setting $L_{D}(a):=\|\partial_{D}(a)\|$. In many interesting cases [AgKa:PSM, Rie:GrpCstar, Christensen-Ivan:spectral-triples], but not always [KyedNest, PutJul:subshifts], $L_{D}$ is a Lip-norm, in which case $({\mathcal{A}},H,D)$ is called a spectral metric space.
To illustrate the complexity of the situation, we now consider the fundamental example, dating back to Connes’ seminal paper [Con:CFH], arising from a countable discrete group $\Gamma$ equipped with a proper length function $\ell\colon\Gamma\rightarrow[0,\infty)$. That is, the function $\ell$ satisfies
@@
From a length function $\ell$, one obtains a spectral triple $({\mathbb{C}}\Gamma,\ell^{2}(\Gamma),D_{\ell})$ on the reduced group $C^{*}$-algebra $C^{*}_{{\operatorname{red}}}(\Gamma)$, by defining $D_{\ell}$ to be the closure of the unbounded operator $\mathcal{D}_{\ell}\colon\text{span}\{\delta_{x}\mid x\in\Gamma\}\rightarrow\ell^{2}(\Gamma)$ defined on the standard orthonormal basis by $\mathcal{D}_{\ell}(\delta_{x}):=\ell(x)\cdot\delta_{x}$. Even in this case, it is very unclear when $L_{D_{\ell}}$ is a Lip-norm — at the time of writing, this is basically only known for groups of polynomial growth [ChristRieffel2017] and hyperbolic groups [OzawaRieffel2005].
@@
Theorem 2.6 ([Rie:MSA, Theorem 1.8]).
The pair $({\mathcal{X}},L)$ is a compact quantum metric space if and only if the set $\{x\in{\mathcal{X}}\mid L(x)\leqslant 1\}$ projects to a totally bounded set in $\mathcal{X}$$\mathbb{C}$} via the quotient map ${\mathcal{X}}\ni x\mapstochar\rightarrow[x]\in{\mathcal{X}}/{\mathbb{C}}$.
@@
[…]
Having finite diameter, in the sense just defined, is equivalent to the (extended) metric $d_{L}$ assigning a finite diameter to the state space ${\mathcal{S}}({\mathcal{X}})$ [Rie:MSA, Proposition 1.6]. Note also, that by connectedness and compactness of ${\mathcal{S}}(\mathcal{X})$ in the weak topology, a compact quantum metric space automatically has finite diameter. Kaad’s criterion now reads as follows:
Theorem 2.8 ([Kaad2023, Theorem 3.1]).
@@
    For every ${\varepsilon}>0$ there exist an operator system ${\mathcal{X}}_{\varepsilon}$ and unital, bounded maps $\iota,\Phi\colon{\mathcal{X}}\rightarrow{\mathcal{X}}_{\varepsilon}$ such that $\iota$ is an isometry, $\Phi$ is positive with finite-dimensional image and $\|\iota(x)-\Phi(x)\|\leqslant{\varepsilon}\cdot L(x)$ for all $x\in{\mathcal{X}}$.